# Device-Independent Certification of Multipartite Distillable Entanglement

### v1

@@ I Introduction
This raises the question: how can we certify or guarantee that a minimum amount of entanglement is produced by a certain procedure or method? This question becomes more interesting when we recall that quantum networks may have components based on different architectures: for example, superconducting qubits for computing [14, 15, 16], trapped ions [17] or solid-state qubits [18] for memory, photons for communication [13], and so on. Hence, we need to adopt a framework of certification that is independent of the constituent parts of the quantum network. This brings us to device-independent (DI) protocols.
The general interest in DI protocols [19, 20] stems from the fact that one can test and verify the amount of entanglement in a state using only classical inputs and outputs, along with classical communication. DI protocols are based on various self-testing results [21, 22, 23, 24], the first of which were developed in [21]. (See [25] for a review.) Most importantly, such protocols are agnostic to the underlying mechanics and specifics of the state being tested. This is why, relevant to the present paper, DI protocols are suitable for the certification of multipartite entanglement in quantum networks, especially ones that contain elements with dissimilar underlying technologies.
There has been a lot of recent interest in device-independent protocols involving more than two parties. This includes investigations into Bell-type inequalities [26, 27] and their application toward DI conference key agreement, including achievable rates [28, 29] and upper bounds on key-agreement rates [30, 31]. However, little attention has been paid toward the device-independent certification of multipartite distillable entanglement. Multipartite entanglement distillation is an important step to ensure that we are using states that are as close as possible to ideal states.
@@
In this work, we extend DI entanglement distillation certification from the bipartite scenario [32] to the multipartite scenario. In what follows, we first give a detailed description and definition of what we mean by DI multipartite entanglement distillation. In Section II, we define a DI, $M$-partite entanglement distillation certification protocol, which consists of completeness and soundness conditions. In Section III, we provide a detailed description of a proposed protocol. Our proposed protocol is centered around the Mermin–Ardehali–Belinskii–Klyshko (MABK) inequality [33, 34, 35], which is critical to proving the completeness condition. We show that the proposed protocol is complete in Section IV. Thereafter, in Section V, we proceed to prove that it is sound, by making use of the entropy accumulation theorem [36] and the structure of the MABK inequality [37].
@@ II Definitions and Setup
The one-shot multipartite distillable entanglement of a state quantifies the amount of multipartite entanglement we can distill from a single copy of the state. It is defined formally as follows:
@@
$E_{D}^{\varepsilon}(\rho_{A_{[M]}})\coloneqq\\ \sup_{\mathcal{L}\in\mathrm{LOCC}}\left\{\log_{2}{d}:F(\mathcal{L}_{A_{[M]}\rightarrow\hat{A}_{[M]}}(\rho),\Phi_{\hat{A}_{[M]}}^{d})\geq 1-\varepsilon\right\},$ (2)
where the optimization is over every LOCC channel, $F(\rho,\sigma)\coloneqq\left\|\sqrt{\rho}\sqrt{\sigma}\right\|_{1}^{2}$, $\Phi_{\hat{A}_{[M]}}^{d}\equiv|\Phi^{d}\rangle\!\langle\Phi^{d}|_{\hat{A}_{[M]}}$ is an $M$-party, rank-$d$ GHZ state, with
@@
$E_{D}(\rho_{A_{[M]}})\coloneqq\inf_{\varepsilon\in(0,1)}\liminf_{n\rightarrow\infty}E_{D}^{n,\varepsilon}(\rho_{A_{[M]}}),$ (4)
where
$E_{D}^{n,\varepsilon}(\rho_{A_{[M]}})\coloneqq\frac{1}{n}E_{D}^{\varepsilon}(\rho_{A_{[M]}}^{\otimes n}).$ (5)
@@
Let $n\in\mathbb{N}$, let $\varepsilon_{\mathrm{dist}}$, $\varepsilon_{\operatorname{snd}}$, $\varepsilon_{\mathrm{cmp}}\in[0,1]$, and let $r\in[0,1]$ be a threshold $M$-partite distillation rate. Furthermore, let $\mathcal{S}^{\mathrm{honest}}$ be a set of “honest” states, each of which is denoted by $\phi^{\mathrm{honest}}$. Let $\mathcal{D}^{\mathrm{honest}}$ be the set of “honest” measurement devices. Let $\mathcal{P}$ be a protocol employing only multipartite LOCC, which, upon being given a state $\sigma\in\mathcal{D}\!\left(\bigotimes_{i=1}^{M}\mathcal{H}_{A_{i}}^{\otimes n}\right)$, creates a state $\rho\in\mathcal{D}\!\left(\bigotimes_{i=1}^{M}\mathcal{H}_{A_{i}}^{\otimes n}\right)$. Let $\rho_{|\Omega}$ denote the final state conditioned on the protocol not aborting.
@@
    Entanglement certification (soundness): For every source $\sigma$ and measurement device, either $E_{D}^{n,\varepsilon_{\operatorname{snd}}}(\rho_{|\Omega})\geq r$ or $\mathcal{P}$ aborts with probability greater than $1-\varepsilon_{\operatorname{snd}}$ when applied on $\sigma$.
To design a protocol and show that it satisfies all the conditions in Definition 3 need to use some information theoretic quantities. Here we recall the definitions of von Neumann entropy, coherent information, and smooth conditional max-entropy.
@@
$I(A\rangle B)_{\rho}\coloneqq H(B)_{\rho}-H(AB)_{\rho}\eqqcolon-H(A|B)_{\rho},$ (7)
where $H(B)_{\rho}$ is the von Neumann entropy of $\rho_{B}=\operatorname{Tr}_{A}[\rho_{AB}]$ and $H(A|B)_{\rho}$ is the quantum conditional entropy of the state $\rho_{AB}$.
@@ III Device-Independent Multipartite Entanglement Certification Protocol
Let $\hat{O}_{0}^{i}$ and $\hat{O}_{1}^{i}$ be two binary observables for all $i\in[M]$. The $M$-partite MABK operator, $\mathcal{K}_{M}$, is defined by the following recursion relation:
@@
The MABK inequalities are such that a violation of the inequalities for $m=1$ proves that at least two parties are entangled, the violation of the inequalities for $m=M-1$ proves genuine $M$-partite entanglement, and the case where $m=M$ gives an upper bound (tight) on what is achievable by quantum mechanics. In this work, we are interested in $m=M-1$ and $m=M$ which correspond to
@@
where $x_{i}\in\{0,1\}$ is the $i$th bit of $x$, and $f:\{0,1\}^{M}\rightarrow\{0,1,\perp\}$ is a function such that $(-1)^{\perp}=0$ by convention.
@@
where $f$ is defined by the $M$-MABK operator in (16). The minimum and maximum winning probabilities when the underlying state is genuinely multipartite entangled are respectively as follows:
@@
Based on the MABK game defined above, we propose Protocol 1. The main result of our paper is that Protocol 1 is a DIMEC protocol, satisfying the requirements of Definition 3, and we state this result formally in Theorem 9.
@@ IV Completeness
In this section, we will show that Protocol 1 is complete. To prove the completeness condition in Definition 3, we need to show that our protocol applied to an arbitrary $\phi^{\operatorname{honest}}\in\mathcal{S}^{\operatorname{honest}}$ will abort with only a small probability. We prove the same in the following theorem.
@@
The probability that Protocol 1 aborts for $\phi^{\operatorname{honest}}\in\mathcal{S}^{\operatorname{honest}}$ is at most
$\varepsilon_{\operatorname{cmp}}\leq\exp\left(-2n\delta^{2}_{\operatorname{est}}\right).$ (24)
Proof. Protocol 1 aborts when $W$, our estimator for the MABK game winning probability, is not high enough. This happens when the MABK violation is not large or if the number of samples is not sufficiently large. Whenever $\phi^{\operatorname{honest}}\in\mathcal{S}^{\operatorname{honest}}$ and the rounds are independent and identically distributed (IID), i.e., an honest implementation, the sequence $(\chi(T_{j}=1)\,W_{j})_{j=1}$ is a sequence of IID random variables. Using Hoeffding’s inequality, we conclude that
$\varepsilon_{\operatorname{cmp}}=\text{Pr}\left(W\leq(\omega_{\mathrm{exp}}\gamma-\delta_{\mathrm{est}})\cdot n\right)\leq e^{-2n\delta^{2}_{\operatorname{est}}},$ (25)
@@
Now, with the completeness of Protocol 1 established, we can proceed to proving its soundness or noise resilience in the following section.
@@ V Soundness
Consider the following setting: a pure state $\psi_{A_{1}\cdots A_{M}BR}$ is held by $M+2$ parties, with the $i$th sender holding $A_{i}$, for $i\in\{1,\ldots,M\}=[M]$, one receiver holding $B$, and a reference system holding $R$, which serves as a purifying system. Additionally, Alice-$i$ and Bob share a maximally entangled state $\Phi_{A^{\prime}_{i}B^{\prime}_{i}}$ of Schmidt rank $c_{i}$, and so the initial overall state is
@@
Multiparty state merging is an LOCC protocol that takes this initial state to
@@
such all of the systems $\hat{A}_{1}\cdots\hat{A}_{M}\hat{B}$ are in Bob’s possession. (For details, see Section 6.3 of [40].) If $c_{i}>d_{i}$, then the protocol consumes $\log_{2}c_{i}-\log_{2}d_{i}$ ebits shared between $A_{i}$ and $B$. If $c_{i}<d_{i}$, then the protocol produces $\log_{2}d_{i}-\log_{2}c_{i}$ ebits, which are shared between $A_{i}$ and $B$ (recall that an ebit is synonymous with the standard Bell pair). If the protocol distills ebits between each Alice, $A_{1}\cdots A_{M}$, and Bob $B$, then they can apply local measurements to convert several copies of their ebits into multipartite GHZ states.
@@
Given the setting above, quantum state merging can be achieved successfully, in the sense that the fidelity constraint in (2) holds, for $\varepsilon=\varepsilon^{\prime}\cdot 4\cdot 3^{-M/2}(2-\varepsilon^{\prime}\cdot 4\cdot 3^{-M/2})\in(0,1)$, if for all $K\subseteq[M]$ such that $K\neq\emptyset$,
@@
Let $\rho_{A_{1}\cdots A_{M}}$ be held by m parties. The following is an achievable rate for GHZ distillation under LOCC:
@@
The statement above gives a bound only on the asymptotic rate of GHZ distillation from mixed states. For our purpose, we need to lower bound $E_{D}^{n,\varepsilon}(\rho_{A_{[M]}})$.
@@
Let $\rho_{A_{1}\cdots A_{M}}$ be held by m parties. The following is an achievable one-shot rate for GHZ distillation under LOCC:
$\max_{k\in[M]}\left\{\min_{K\subseteq[M]\backslash k}\frac{-H^{\varepsilon^{\prime}}_{\max}(A_{K}|A_{[M]\backslash K})_{\rho}}{|K|}\right\}+\frac{2\log\varepsilon^{\prime}}{M-1}.$ (30)
Proof. To achieve this one-shot rate , we can use the protocol outlined in [41, Theorem 9] but without time sharing. Time shared decoding is replaced with simultaneous decoding, as the latter is necessary for the one-shot setting. We obtain achievable rates for the one-shot case by replacing coherent information in (29) with the conditional smooth max-entropy. This leaves us with an expression commensurate with (28) in Theorem 2.
V.1 Entropy Accumulation Theorem
@@
For a sequential process like our protocol, the most restrictive relation between the rounds is to assume that rounds are IID. However, such a restriction is arguably too strong for the device-independent scenario. Entropy accumulation theory allows for a more general relation between the rounds of our protocol. This is encapsulated in the definition of EAT channels. EAT channels are defined as follows.
@@
If rounds are not IID then, we cannot consider a round of the protocol to be completely independent of the preceding rounds. In other words, we cannot use the additivity property of the quantities under consideration. To resolve this, entropy accumulation theory requires that a special function, associated with our protocol, be found. This special function is called the max-tradeoff function.
To define max-tradeoff functions, we need to clarify some notation. We will use the following notation henceforth. Given a value $w=(w_{1},\dotsc,w_{n})\in\mathcal{W}^{n}$, where $\mathcal{W}$ is a finite alphabet, we denote by $\mathrm{freq}_{w}$ the probability distribution over $\mathcal{W}$ defined by $\mathrm{freq}_{w}(\tilde{w})\coloneqq\frac{|\left\{j|w_{j}=\tilde{w}\right\}|}{n}$ for $\tilde{w}\in\mathcal{W}$. If $\tau$ is a state classical on $W$, we write $\Pr\left[w\right]_{\tau}$ to denote the probability that $\tau$ assigns to $w$. Now, we move on to the defintion of max-tradeoff functions.
@@
The statement of the EAT, relevant for the smooth max-entropy, is given below (see [36, Theorem 4.4] and Eq. (A.2) of [43, Appendix A]).
@@
$v\coloneqq 2\left(\log_{2}(1+2d_{O_{i}})+\lceil\left\|\triangledown f_{\max}\right\|_{\infty}\rceil\right)\\ \qquad\times\sqrt{1-2\log_{2}(\varepsilon\cdot\Pr[\Omega]_{\tau})}.$ (33)
It has been shown that Theorem 4, the generalized entropy accumulation theorem, holds regardless of whether the sequence of channels $\{\mathcal{N}_{j}\}_{j}$ satisfies a Markov-chain condition [43, Appendix A]. Note that the original entropy accumulation theorem [36] does require such a Markov-chain condition.

### v2

@@ I Introduction
This point raises the question: how can we certify or guarantee that a minimum amount of entanglement is produced by a certain procedure or method? This question becomes more interesting when we recall that quantum networks may have components based on different architectures: for example, superconducting qubits for computing [14, 15, 16], trapped ions [17] or solid-state qubits [18] for memory, photons for communication [13], and so on. Hence, we need to adopt a framework of certification that is independent of the constituent parts of the quantum network, thus bringing us to device-independent (DI) protocols.
The general interest in DI protocols [19, 20] stems from the fact that one can test and verify the amount of entanglement in a state using only classical inputs and outputs, along with classical communication. DI protocols are based on various self-testing results [21, 22, 23, 24], the first of which were developed in [21]. (See [25] for a review.) Most importantly, such protocols are agnostic to the underlying mechanics and specifics of the state being tested. As such, and relevant to the present paper, DI protocols are suitable for the certification of multipartite entanglement in quantum networks, especially ones that contain elements with dissimilar underlying technologies.
There has been a lot of recent interest in device-independent protocols involving more than two parties, including investigations into Bell-type inequalities [26, 27] and their application toward DI conference key agreement, involving achievable rates [28, 29] and upper bounds on key-agreement rates [30, 31]. However, little attention has been paid toward the device-independent certification of multipartite distillable entanglement. Multipartite entanglement distillation is an important step to ensure that we are using states that are as close as possible to ideal states.
@@
In more detail, we extend DI entanglement distillation certification from the bipartite scenario [32] to the multipartite scenario. In what follows, we first give a detailed description and definition of what we mean by DI multipartite entanglement distillation. In Section II, we define a DI, $M$-partite entanglement distillation certification protocol, which consists of completeness and soundness conditions. In Section III, we provide a detailed description of a proposed protocol. Our proposed protocol is centered around the Mermin–Ardehali–Belinskii–Klyshko (MABK) inequality [33, 34, 35], which is critical to proving the completeness condition. We show in Section IV that the proposed protocol is complete. Thereafter, in Section V, we proceed to proving that it is sound, by making use of the entropy accumulation theorem [36] and the structure of the MABK inequality [37].
@@ II Definitions and Setup
In this section, we provide several definitions that we use throughout the rest of the paper, including the GHZ orthonormal basis, the one-shot multipartite distillable entanglement of a multipartite quantum state, an entanglement distillation certification protocol, and basic entropies.
@@
The one-shot multipartite distillable entanglement of a state quantifies the amount of multipartite entanglement that we can distill from a single copy of the state. It is defined formally as follows:
@@
$E_{D}^{\varepsilon}(\rho_{A_{[M]}})\coloneqq\\ \sup_{\begin{subarray}{c}d\in\mathbb{N},\\ \mathcal{L}\in\mathrm{LOCC}\end{subarray}}\left\{\log_{2}{d}:F(\mathcal{L}_{A_{[M]}\rightarrow\hat{A}_{[M]}}(\rho),\Phi_{\hat{A}_{[M]}}^{d})\geq 1-\varepsilon\right\},$ (2)
where the optimization is over every LOCC channel $\mathcal{L}_{A_{[M]}\rightarrow\hat{A}_{[M]}}$, the fidelity is defined as $F(\rho,\sigma)\coloneqq\left\|\sqrt{\rho}\sqrt{\sigma}\right\|_{1}^{2}$, and $\Phi_{\hat{A}_{[M]}}^{d}\equiv|\Phi^{d}\rangle\!\langle\Phi^{d}|_{\hat{A}_{[M]}}$ is an $M$-party, rank-$d$ GHZ state, with
@@
$E_{D}(\rho_{A_{[M]}})\coloneqq\inf_{\varepsilon\in(0,1)}\liminf_{n\rightarrow\infty}\frac{1}{n}E_{D}^{\varepsilon}(\rho_{A_{[M]}}^{\otimes n}).$ (4)
@@
Let $n\in\mathbb{N}$, let $\varepsilon_{\mathrm{smo}}$, $\varepsilon_{\operatorname{snd}}$, $\varepsilon_{\mathrm{cmp}}\in[0,1]$, and let $r\geq 0$ be a threshold $M$-partite distillation rate. Furthermore, let $\mathcal{S}^{\mathrm{honest}}$ be a set of “honest” states, each of which is denoted by $\phi^{\mathrm{honest}}$. Let $\mathcal{D}^{\mathrm{honest}}$ be the set of “honest” measurement devices. Let $\mathcal{P}$ be a protocol employing only multipartite LOCC, which, upon being given a state $\sigma\in\mathcal{D}\!\left(\bigotimes_{i=1}^{M}\mathcal{H}_{A_{i}}^{\otimes n}\right)$, creates a state $\rho\in\mathcal{D}\!\left(\bigotimes_{i=1}^{M}\mathcal{H}_{A_{i}}^{\otimes n}\right)$. Let $\rho_{|\Omega}$ denote the final state conditioned on the protocol not aborting.
@@
    Entanglement certification (soundness): For every source $\sigma$ and measurement device, either $E_{D}^{n,\varepsilon_{\operatorname{snd}}}(\rho_{|\Omega})\geq r$ or $\mathcal{P}$ aborts with probability greater than $1-\varepsilon_{\operatorname{smo}}$ when applied on $\sigma$.
To design a protocol and show that it satisfies all the conditions in Definition 3, we need to use some information-theoretic quantities. Here we recall the definitions of von Neumann entropy, coherent information, and smooth conditional max-entropy.
@@
$I(A\rangle B)_{\rho}\coloneqq H(B)_{\rho}-H(AB)_{\rho}=-H(A|B)_{\rho},$ (6)
where $H(B)_{\rho}$ is the von Neumann entropy of $\rho_{B}=\operatorname{Tr}_{A}[\rho_{AB}]$ and $H(A|B)_{\rho}\coloneqq H(AB)_{\rho}-H(B)_{\rho}$ is the quantum conditional entropy of the state $\rho_{AB}$.
@@ III Device-Independent Multipartite Entanglement Certification Protocol
Let $\hat{O}_{0}^{i}$ and $\hat{O}_{1}^{i}$ be binary observables for all $i\in[M]$. The $M$-partite MABK operator, $\mathcal{K}_{M}$, is defined by the following recursion relation:
@@
The MABK inequalities are such that a violation of the inequalities for $m=1$ proves that at least two parties are entangled, the violation of the inequalities for $m=M-1$ proves genuine $M$-partite entanglement, and the case where $m=M$ gives an upper bound (tight) on what is achievable by quantum mechanics. In this work, we are interested in $m=M-1$ and $m=M$, which correspond to
@@
where $x_{i}\in\{0,1\}$ is the $i$th bit of $x$ and $f:\left\{0,1\right\}^{M}\rightarrow\left\{0,1,\perp\right\}$ is a function such that $(-1)^{\perp}=0$ by convention.
@@
where $f$ is defined by the $M$-MABK operator in (15). The minimum and maximum winning probabilities when the underlying state is genuinely multipartite entangled are respectively as follows [37]:
@@
Based on the MABK game defined above, we propose Protocol 1. There are two types of rounds in Protocol 1: a testing round, where the parties play the MABK game and record the result, and a storage round, where the parties simply store the state they receive in a quantum memory. At the beginning of each round, we choose between these types uniformly random. The main result of our paper is that Protocol 1 is a DIMEC protocol, satisfying the requirements of Definition 3, and we state this result formally in Theorem 9.
@@ IV Completeness
In this section, we show that Protocol 1 is complete. To prove the completeness condition in Definition 3, we need to show that our protocol applied to an arbitrary $\phi^{\operatorname{honest}}\in\mathcal{S}^{\operatorname{honest}}$ aborts with only a small probability. We prove this result in the following theorem.
@@
The following bound holds:
$\varepsilon_{\operatorname{cmp}}\leq\exp\!\left(-2n\delta^{2}_{\operatorname{est}}\right),$ (23)
implying that the probability that Protocol 1 aborts for $\phi^{\operatorname{honest}}\in\mathcal{S}^{\operatorname{honest}}$ is no larger than $\exp\!\left(-2n\delta^{2}_{\operatorname{est}}\right)$.
Proof. Protocol 1 aborts when $W$, the estimator for the MABK game winning probability, is not high enough. This happens when the MABK violation is not large or if the number of samples is not sufficiently large. Whenever $\phi^{\operatorname{honest}}\in\mathcal{S}^{\operatorname{honest}}$ and the rounds are independent and identically distributed (IID), i.e., an honest implementation, the sequence $\left(\chi(T_{j}=1)\,W_{j}\right)_{j=1}$ is a sequence of IID random variables. Using Hoeffding’s inequality [40], we conclude that
$\varepsilon_{\operatorname{cmp}}=\Pr\!\left(W\leq(\omega_{\mathrm{exp}}\gamma-\delta_{\mathrm{est}})\cdot n\right)\leq e^{-2n\delta^{2}_{\operatorname{est}}},$ (24)
@@ V Soundness
Consider the following setting: a pure state $\psi_{A_{1}\cdots A_{M}BR}$ is held by $M+2$ parties, with the $i$th sender holding $A_{i}$, for $i\in\{1,\ldots,M\}\eqqcolon[M]$, one receiver holding $B$, and a reference system holding $R$, which serves as a purifying system. Additionally, Alice-$i$ and Bob share a maximally entangled state $\Phi(c_{i})_{A^{\prime}_{i}B^{\prime}_{i}}$ of Schmidt rank $c_{i}\in\mathbb{N}$, and so the initial overall state is
@@
Multiparty state merging is an LOCC protocol that transforms this initial state to
@@
such all of the systems $\hat{A}_{1}\cdots\hat{A}_{M}\hat{B}$ are in Bob’s possession. (For details, see Section 6.3 of [41].) If $c_{i}>d_{i}$, then the protocol consumes $\log_{2}c_{i}-\log_{2}d_{i}$ ebits shared between $A_{i}$ and $B$. If $c_{i}<d_{i}$, then the protocol produces $\log_{2}d_{i}-\log_{2}c_{i}$ ebits, which are shared between $A_{i}$ and $B$ (recall that the term “ebit” is synonymous with the standard Bell pair). If the protocol distills ebits between each Alice, $A_{1}\cdots A_{M}$, and Bob $B$, then they can apply LOCC to convert several copies of their ebits into multipartite GHZ states.
@@
Given the setting above, quantum state merging can be achieved successfully, in the sense that the fidelity constraint in (2) holds, for
$\varepsilon=8\cdot 3^{M/2}\sqrt{\varepsilon^{\prime}}\in(0,1),$ (27)
if for all $K\subseteq[M]$ such that $K\neq\emptyset$,
@@
Let $\rho_{A_{1}\cdots A_{M}}$ be a state held by $m$ parties. The following is an achievable rate for GHZ distillation under LOCC:
@@
The statement above gives a bound only on the asymptotic rate of GHZ distillation from mixed states. For our purpose, we require a lower bound on $\frac{1}{n}E_{D}^{\varepsilon}(\rho_{A_{[M]}}^{\otimes n})$. The following proposition gives such a bound by making use of the developments of [41, 42].
@@
Let $\rho_{A_{1}\cdots A_{M}}$ be held by $m$ parties. Fix $\varepsilon^{\prime}>0$ such that $\varepsilon\coloneqq 8\cdot 3^{M/2}\sqrt{\varepsilon^{\prime}}\in(0,1)$. Then the following is an achievable one-shot rate for GHZ distillation under LOCC; i.e.,
$\frac{1}{n}E_{D}^{\varepsilon}(\rho_{A_{[M]}}^{\otimes n})\geq\\ \max_{k\in[M]}\left\{\min_{K\subseteq[M]\backslash k}\frac{-H^{\varepsilon^{\prime}}_{\max}(A_{K}|A_{[M]\backslash K})_{\rho}}{|K|}\right\}+\frac{2\log\varepsilon^{\prime}}{M-1}.$ (30)
Proof. To achieve this one-shot rate, we can use the protocol outlined in [42, Theorem 9] but without time sharing. Time shared decoding is replaced with simultaneous decoding, as the former is not possible in the one-shot setting. We obtain achievable rates for the one-shot case by replacing coherent information in (29) with the conditional smooth max-entropy. This leaves us with an expression commensurate with (28) in Theorem 2.
V.1 Entropy Accumulation Theorem (EAT)
@@
For a sequential process like our protocol, the most restrictive relation between the rounds is to assume that they are IID. However, such a restriction is arguably too strong for the device-independent scenario. Entropy accumulation theory allows for a more general relation between the rounds of our protocol. This is encapsulated in the following definition of EAT channels.
@@
If the rounds are not IID, then we cannot consider a round of the protocol to be completely independent of the preceding rounds. In other words, we cannot use the additivity property of the quantities under consideration. To resolve this, entropy accumulation theory requires that a special function, associated with our protocol, be found. This special function is called the max-tradeoff function.
To define max-tradeoff functions, we need to clarify some notation, which we will use henceforth. Given a value $w=(w_{1},\dotsc,w_{n})\in\mathcal{W}^{n}$, where $\mathcal{W}$ is a finite alphabet, we denote by $\mathrm{freq}_{w}$ the probability distribution over $\mathcal{W}$ defined by $\mathrm{freq}_{w}(\tilde{w})\coloneqq\frac{\left|\left\{j|w_{j}=\tilde{w}\right\}\right|}{n}$ for $\tilde{w}\in\mathcal{W}$. If $\tau$ is a state classical on $W$, we write $\Pr\left[w\right]_{\tau}$ to denote the probability that $\tau$ assigns to $w$. Now, we move on to the definition of max-tradeoff functions.
@@
The statement of the EAT, relevant for the smooth max-entropy, is given below (see [36, Theorem 4.4] and Eq. (A.2) of [44, Appendix A]).
@@
$v\coloneqq 2\left(\log_{2}(1+2d_{O_{i}})+\lceil\left\|\triangledown f_{\max}\right\|_{\infty}\rceil\right)\\ \qquad\times\sqrt{1-2\log_{2}(\varepsilon_{\mathrm{smo}}\cdot\Pr[\Omega]_{\tau})}.$ (33)
It has been shown that Theorem 4, the generalized entropy accumulation theorem, holds regardless of whether the sequence of channels $\{\mathcal{N}_{j}\}_{j}$ satisfies a Markov-chain condition [44, Appendix A]. Note that the original entropy accumulation theorem [36] does require such a Markov-chain condition.